Circular No. 6158

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995J IN NGC 4512 A. V. Filippenko, University of California at Berkeley, reports that

T. Hurt (University of California, Santa Barbara) obtained CCD spectra (range 340–800 nm) of SN 1995J (cf. IAUC 6157) on Apr. 9 UT with the 3m Shane reflector at Lick Observatory. The partially-calibrated data reveal very strong, broad P-Cyg profiles of hydrogen, indicating that this is a normal type-II supernova ≲ 2 months past maximum brightness.

S. Nakano, Sumoto, Japan, forwards the following precise position measured by Y. Kushida from a CCD frame taken by R. Kushida at Yatsugatake South Base Observatory on Apr. 7.659 UT: α = 12^{h}32^{m}38^{s}.34, δ = +63^{o}53′09′′.2 (equinox 2000.0); this position corresponds to a point 29′′ east and 30′′ north of the galaxy’s center. The unfiltered magnitude of SN 1995J was found to be 16.1, using V GSC comparison-star magnitudes. A measurement obtained at Sormano from a CCD observation by P. Sicoli and V. Giuliani on Apr. 7.944 UT gives position end figures 38^{s}.31, 09′′.4.

SUPERNOVA 1995I IN ANONYMOUS GALAXY P. Garnavich, P. Challis, and R. Kirshner, Center for Astrophysics,

report that spectra obtained with the Multiple Mirror Telescope on Mar. 31 indicate that SN 1995I (cf. IAUC 6153) is a type-II supernova. A broad emission feature with a weak P-Cyg profile has developed at the rest wavelength of Hα. Spectra taken on Mar. 9 at the MMT showed only a blue continuum that peaked at a rest wavelength of 400 nm. The redshift is 0.075, based on narrow emission lines from the host galaxy.

V1974 CYGNI A. Retter, E. O. Ofek, and E. M. Leibowitz, Wise Observatory, Tel

Aviv University, report: “Photometry of V1974 Cyg (N Cyg 1992) in the I band during seven nights in 1994 confirms the finding of Semeniuk et al. (1994, Acta Astron. 44, 277) about the presence of a periodicity of 0.08505 day in the light curve of the star. This period is 4.7 percent longer than a second period of 0.08127 day, which is also clearly present in the light curve and which was identified in 1993 as the binary periodicity of the system by DeYoung and Schmidt (1994, Ap.J. 431, L47). The two periods obey exactly the linear relation of Stolz and Schoembs (1981, IBVS 2029) for SU UMa stars.”

1995 April 10 Daniel W. E. Green